OpenStudy (anonymous):

What are the posible values of x such that 3x^2 - 2x = 0

OpenStudy (anonymous):

answers are: -2/3 only, 0 only, 2/3 only, 0 and 2/3

OpenStudy (anonymous):

Since x is a common factor in both terms on the left side of the equations, you can factor out an x. x(3x-2)=0 Then solve each term by setting equal to zero. x=0 3x-2=0 3x=2 Subtract from both sides. x=2/3 Divide both sides by 3 to isolate x by itself. Answer is values for x are 0 and 2/3

OpenStudy (anonymous):

you can factor: \[3x^2-2x=0\rightarrow x(3x-2)=0\] 0 and 2/3

OpenStudy (anonymous):

values for x are 0 and 2/3

OpenStudy (anonymous):

Yess sir..thanks guys